A pole reported that 64% of adults were satisfied with the job the major airlines were doing. Suppose 25 adults were selected at random and the number who are satisfied is recorded question E would it be unusual to find more than 22 who are satisfied with the job and major airlines were doing.
the question is:
The result (is or is not) unusual because P(x>22)=(fill in the blank) under the assumption that the proportion of adults that are satisfied with the airline is 64%. Thus, in 100 random samples of size 25 this result is expected in about (fill in the blank) of the random samples.
Answer #1

We need to find the P(x>22)

P(X>22) = 1- (P(X=23)+P(X=24)+P(X=25))

P(X=23)

P(X=24)


P(X=25)

P(X>22) = (P(X=23)+P(X=24)+P(X=25))
(X>22) = (0.0014 + 0.0002+ 0)=0.00015

The result (is ) unusual because P(x>22)=(0.00015) under the assumption that the proportion of adults that are satisfied with the airline is 64%.

Thus, in 100 random samples of size 25 this result is expected in about 16 of the random samples.